An unknown salt is added continously to water until salt is no longer dissolved. Initially, the water weighs 50 g. When the mixture reached 65 g, salt can no longer be dissoleved. What is the saturation concentration of the salt in g/L?
1
Expert's answer
2020-10-12T13:30:17-0400

m(salt) = 65 - 50 = 15 g

saturation concentration = 15 / 50 * 1000 = 300 g/L